# AUDETER: A Large-scale Dataset for Deepfake Audio Detection in Open Worlds

### v1

@@ 4.3.3. Effect of Synthetic Pattern Diversity in Large-scale Training
CV Val CV Test PS Test Single Sys. TTS AVG Voc Avg TTS AVG Voc Avg TTS AVG Voc Avg SparkTTS 0.096 0.496 0.100 0.496 0.482 0.541 F5-TTS 0.150 0.492 0.155 0.491 0.660 0.569 Fish-Speech 0.093 0.438 0.097 0.442 0.510 0.464 XTTS 0.172 0.496 0.178 0.496 0.669 0.531 VITS 0.108 0.485 0.111 0.485 0.455 0.533 BigVGAN 0.547 0.270 0.565 0.296 0.495 0.426 HiFi-GAN 0.238 0.285 0.249 0.290 0.691 0.441 Vocos 0.262 0.062 0.302 0.090 0.613 0.320 UnivNet 0.233 0.285 0.212 0.126 0.399 0.227 Mel GAN 0.235 0.199 0.245 0.201 0.697 0.315
@@ 5. Conclusion and Future Works
In this paper, we introduce AUDETER, a large-scale deepfake audio detection dataset for systematic benchmarking and robust large-scale training of deepfake audio detectors in open-world detection. AUDETER contains nearly 3 million synthetic audio samples generated by 21 recent speech synthesis systems that correspond to diverse selections of real human speech from 4 corpora, enabling fine-grained and balanced evaluation across human voice domains and speech systems, under various domain shifts. Through extensive experiments using the complete dataset, we demonstrate that AUDETER effectively challenges existing detection models and reveals their limitations. We also conduct large-scale training experiments using different subset combinations of AUDETER. These experiments achieve superior open-world detection performance, demonstrating AUDETER’s value as a resource for data-centric improvements and highlighting the importance of high-quality training data. We show that using high-volume diverse deepfake samples with large audio backbones effectively improves generalisation to novel audio samples.
We plan to continue developing AUDETER as an ongoing project to address the rapidly evolving nature of speech synthesis systems. We recognise that even large-scale approaches will eventually become outdated as new synthesis methods emerge. Looking forward, we plan to identify and extract representative synthesis patterns that can generalise across multiple systems to avoid continuously scaling up our dataset. Another promising direction is to explore advanced training methodologies for improved generalisation performance, such as self-supervised pretraining.
@@ Appendix A Detailed Dataset Information
A.1. Audio Synthesis Models
@@ A.1.1. TTS Models Selection
A.1.2. Vocoder Models Selection
@@
A.2. Real Audio
We summarise the four corpora of of real huamn voice in Table 11.
@@ A.2. Real Audio
| The People’s Speech | Hugging Face Datasets: MLCommons/peoples_speech (validation and test partition from the clean subset) |
@@
A.3. Detailed Dataset Statistics
@@
| In-the-Wild | Common Voice | People Speech | MLS | |||||||
| Model | Bona-fide | Val | Test | Total | Val | Test | Total | Dev | Test | Total |
@@
| Vocoder Collection | ||||||||||
| In-the-Wild | Common Voice | People Speech | MLS | |||||||
@@ Appendix B Complete Datasets Quality Assurance Results
B.1. Detailed Naturalness Scores
We provide the details results for the NISQA naturalness assessmenton all our deepfake audio samples and present the results in Table 16.
| itw | cv | ps | mls | ||||
@@ B.1. Detailed Naturalness Scores
Appendix C Complete Experimental Results
C.1. Detailed Zero-shot Performance
We present selected results of open-world detection performance of the baseline methods on the Common Voice and Peoples Speech subsets for both the TTS and Vocoder collections due to space limitations. Here, we provide the detailed results for all subsets on both collections.
@@ C.1. Detailed Zero-shot Performance
C.2. Single System Generalisation Result
@@
C.2.1. Same textual information, same domain.
@@
C.2.2. Same domain different textual information.
@@
C.2.3. Different domain and different textual information.

### v2

@@ 4.3.4. Effect of Synthetic Pattern Diversity in Large-scale Training
CS Train CS Test UC Test Single Sys. TTS AVG Voc Avg TTS AVG Voc Avg TTS AVG Voc Avg SparkTTS 9.6 49.6 10.0 49.6 48.2 54.1 F5-TTS 15.0 49.2 15.5 49.1 66.0 56.9 Fish-Speech 9.3 43.8 9.7 44.2 51.0 46.4 XTTS 17.2 49.6 17.8 49.6 66.9 53.1 VITS 10.8 48.5 11.1 48.5 45.5 53.3 BigVGAN 54.7 27.0 56.5 29.6 49.5 42.6 HiFi-GAN 23.8 28.5 24.9 29.0 69.1 44.1 Vocos 26.2 6.2 30.2 9.0 61.3 32.0 UnivNet 23.3 28.5 21.2 12.6 39.9 22.7 Mel GAN 23.5 19.9 24.5 20.1 69.7 31.5
4.4. Towards Diverse Deepfake Training
4.4.1. Limitations of Binary Classification for Training Detection Models
Regardless of model architecture, most existing DFAD models are trained as a binary classification problem. Beyond limited generalisation to unseen systems, this formulation can induce negative transfer when trained on highly diverse deepfake sources. In particular, certain speech synthesis models exhibit strong system specific fingerprints, which can bias the training signal and encourage models to overfit to these cues, limiting the learning of representations that generalise across most systems. We refer to such systems as harmful systems. We observe that including such harmful systems under current training schemes can substantially degrade cross domain performance. To better study this effect, we incorporate two recent evaluation focused datasets: DFADD and SONAR, alongside ITW to strengthen cross-domain evaluation.
4.4.2. Proposed Baseline Method
As a complementary contribution, we propose a two stage curriculum training strategy to mitigate negative transfer caused by harmful systems (Fig. 6). The core intuition is to first establish system-invariant and universal representations using data samples generated by systems that exhibit weaker system specific artefacts, before gradually incorporating those harmful system with more pronounced fingerprints.
The first stage involves two main tasks: identifying harmful systems via linear probing and training a base detector using system that are considered as non-harmful, which we refer to as core-systems. The linear probing test aims to identify systems that expose strong and persistent system specific cues in the representations learned for DFAD. We freeze the backbone of a detector trained on all systems and optimise a lightweight system identification head as a proxy for system harmfulness. After light tuning of the head, harmful systems exhibit balanced precision and recall, indicating persistent system-specific cues that signal early dominance. Figure 7 shows the results of the probing test on AUDETER systems. For our baseline model, we treat the most dominant systems from each subset: SparkTTS for TTS and BigVGAN for vocoders, as harmful systems and use the remaining systems to train a base detector.
In the second stage, we adopt a student–teacher paradigm as a regulariser for supervised training on the full system set. The motivation is to prevent harmful systems from overshadowing the system-invariant cues learned in Stage 1. Strongly dominant systems can exert a disproportionate influence during training, and even when they constitute only a small fraction of the data, they may bias the representation toward system-specific artefacts rather than generalisable cues. Let and denote the teacher (Stage-1) and student models, respectively, producing logits for an input sample ; denote the core system set and the full training set, then in Stage 2, we fix and optimise by minimising
where if and otherwise. The distillation loss is defined as:
This encourages the student model to preserve the generalisable cues learned by the teacher while adapting to newly introduced systems. To further improve training stability, we employ LoRA adapters in the last three layers and freeze all other parameters, unfreezing only the aggregator module to learn multi-layer representations. During inference, the student model is used end-to-end with the LoRA weights merged for scoring, introducing negligible overhead to the base architecture. Framework details are presented in Section C.2.
4.4.3. Cross-dataset Performance
Cross Domain In Domain Overall Model ITW DFAFF SONAR Avg AUDETER Avg RawNet2 37.81 41.86 43.00 40.89 44.20 32.80 AASIST 43.02 31.86 57.47 44.12 51.90 45.67 RawGAT-ST 52.60 23.70 50.78 42.36 23.70 38.63 PC-Dart 66.17 85.17 51.90 67.75 38.40 60.41 SAMO 37.09 78.68 42.08 52.62 59.85 54.43 NVA 26.32 51.97 74.61 50.97 55.90 52.20 Purdue-M2 79.57 58.41 73.00 70.33 44.90 63.97 XLR+R+A 10.46 11.93 46.12 22.84 23.70 23.05 XLR+SLS 7.46 7.54 24.72 13.24 35.80 18.88 XLR+R+A (sup) 2.03 27.68 13.83 14.51 0.87 11.10 XLR+SLS (sup) 1.87 22.80 14.58 13.08 0.71 9.99 XLR+R+A (cur) 2.31 6.27 12.78 7.12 1.48 5.71
Table 9 summarises the results on the extended benchmark. The proposed method achieves substantially lower EER than both pretrained baselines and AUDETER-trained XLR models using all systems in terms of cross-domain and overall performance. Although a minor in-domain performance drop is observed, this trade-off is worthwhile, as the method avoids overfitting to systems with strong synthesis fingerprints and improves cross-domain generalisation. This demonstrates that the proposed method effectively learns balanced and robust representations, avoiding dominance by any single synthesis system.
@@ 5. Conclusion and Future Works
This paper presents AUDETER, a large-scale deepfake audio detection dataset for enhanced evaluation and training generalised detection models for open-world detection. It contains nearly 3 million synthetic audio generated by us using 21 recent speech synthesis systems paired with diverse real speech from 4 corpora, enabling controlled, fine-grained evaluation under various domain shifts. Extensive empirical evaluation shows that AUDETER effectively challenges existing detection models, and that training on AUDETER’s large scale and diverse synthetic patterns leads to significantly improved generalisation. As a complementary contribution, we identify that the default training schemes used by existing methods suffer from negative transfer when trained on diverse deepfake data, and we propose a curriculum learning strategy to mitigate this effect. Looking forward, we view AUDETER as a foundation for developing more effective training paradigms for deepfake detection. In particular, future work will explore self supervised and representation learning approaches tailored to deepfake audio, with the goal of learning robust, system invariant features from diverse training data.
@@ Appendix A Detailed Dataset Information
A.1. Dataset Generation Pipeline
A.2. Audio Synthesis Models
@@ A.2.1. TTS Models Selection
A.2.2. Vocoder Models Selection
@@
A.3. Real Audio
We summarise the four corpora of of real huamn voice in Table 12. The real-speech subsets in AUDETER are named based on their speech characteristics and recording conditions rather than the original dataset titles. Each subset is derived from an existing real-speech corpus and used solely as a source of human speech for synthetic audio generation and evaluation:
Celebrity is derived from the In-the-Wild dataset and contains recordings of public figures captured under diverse real-world conditions.
CrowdSource is derived from Common Voice 13.0, consisting of speech uploaded by public users, with read scripts recorded across a wide range of accents, devices, and acoustic environments.
US Congress is derived from The People’s Speech, which includes formal recordings of U.S. congressional sessions and related public-domain speech.
Audiobook is derived from Multilingual LibriSpeech (MLS) and contains clean, professionally recorded read speech from audiobooks.
@@ A.3. Real Audio
| The People’s Speech | Hugging Face Datasets: MLCommons/peoples_speech (train and test partition from the clean subset) |
@@
A.4. Deepfake Audio
All AUDETER subsets include predefined train and test partitions, with the exception of Celebrity, which is constructed from our curated version of In-the-Wild data and is primarily used for cross-domain evaluation against the standard ITW benchmark.
For within-AUDETER experiments, models are trained on the train split and evaluated on the test split of each subset.
For cross-domain evaluation (e.g., evaluation on ITW, DFADD, or SONAR), both train and test splits of AUDETER subsets may be used for training to provide the strongest possible training coverage.
The Celebrity subset is not used for within-dataset training and testing splits. Instead, it serves as a controlled AUDETER counterpart to the classic ITW dataset for cross-domain comparison.
A.5. Detailed Dataset Statistics
@@
| Celebrity | CrowdSource | US Congress | Audiobook | |||||||
| Model | Test | Val | Test | Total | Val | Test | Total | Test | Test | Total |
@@
| TTS Collection | ||||||||||
| Celebrity | CrowdSource | US Congress | Audiobook | |||||||
@@ Appendix B Complete Datasets Quality Assurance Results
B.1. Detailed Intelligibility Scores
B.2. Detailed Naturalness Scores
We provide the detailed results for the NISQA naturalness assessmenton all our deepfake audio samples and present the results in Table 17.
| Celebrity | CrowdSource | US Congress | Audiobook | ||||
@@ B.2. Detailed Naturalness Scores
Appendix C Implementation Details
C.1. EER Evaluation with Multiple Synthetic Variants
Pseudocode for the evaluation protocol with multiple synthetic variants is provided in Algorithm 1,
C.2. Curriculum Learning
We adopt a two-stage curriculum learning strategy to incorporate unseen TTS systems while preserving performance on previously learned core systems. Stage 1 trains the detector on 19 core TTS methods using standard cross-entropy with RawBoost augmentation. Stage 2 introduces two new systems (SparkTTS and BigVGAN) via teacher–student distillation with LoRA, where rank-4 adapters () are inserted into the Q, K, V, and output projections of XLS-R transformer layers 21–23, with Kaiming-initialised A and zero-initialised B to ensure identity at initialisation. The XLS-R backbone is frozen except for LoRA parameters, while the AASIST graph body remains trainable with a lower learning rate (); LoRA adapters and the classifier use . Training minimises a combined weighted BCE loss over all samples and a KL distillation term applied only to core samples against a frozen Stage-1 teacher (, ). Models are trained for 20 epochs with batch size 128 using Adam, balanced real/fake sampling, and gradient clipping at 1.0 on CrowdSource and US Congress with RawBoost augmentation. Both LoRA-only and merged checkpoints are saved for adapter-free inference, with further details provided in the appendix.
Appendix D Complete Experimental Results
D.1. Effect of Speakers on Detectability.
Table 18 reports the average detection performance of baseline methods across multiple speaker variants of the same synthesis system (OpenVoice V2). We observe substantial performance variation across speakers for all methods, indicating that speaker characteristics alone can significantly affect deepfake detectability, even when the synthesis architecture is fixed. This suggests that system-level evaluation alone is insufficient, as intra-system speaker variability introduces additional domain shifts. While stronger models such as XLR-based detectors show more stable performance across speakers, no method is fully robust to speaker-induced variation. These results highlight the importance of evaluating deepfake detectors under multi-speaker settings and motivate the inclusion of speaker diversity when assessing open-world generalisation.
D.2. Detailed Zero-shot Performance
We present selected results of open-world detection performance of the baseline methods on the Crowdsource and UC Congress subsets for both the TTS and Vocoder collections due to space limitations. Here, we provide the detailed results for all subsets on both collections.
@@ D.2. Detailed Zero-shot Performance
D.3. Single System Generalisation Result
@@
D.3.1. Same textual information, same domain.
@@
D.3.2. Same domain different textual information.
@@
D.3.3. Different domain and different textual information.